# Non-uniqueness of the steady state for run-and-tumble particles with a double-well interaction potential

### v2

@@ I.1 Overview
In this paper we consider one-dimensional RTPs interacting via a pairwise polynomial potential of the form $W(r)=-k_{0}\,r^{2}/2+g\,r^{4}/4$, which for $k_{0},g>0$ is repulsive at short distance and attractive at large distance. Although the interaction force $W^{\prime}(r)$ diverges at large distance, which is a priori an unrealistic feature (see however below), it has two advantages from a computational perspective. First, $W^{\prime}(0)=0$, hence it allows for particle crossings and therefore allows to use the extension of the Dean-Kawasaki method Dean ; Kawa to the RTP (while it was shown that this method fails to describe single-file active particles TouzoDBM2023 ). Second, it is simple enough so that the self-consistent equation obeyed by the stationary density in the large $N$ limit, derived in activeRDshort , can be solved analytically. We find that in the steady state the system forms a bound state with a bounded support, which exhibits a transition between a joint support and a disconnected support where particles spontaneously split into two groups. Furthermore, remarkably, we find that the steady state is not unique, and that in some range of parameters it exhibits bistability, as well as a breaking of the parity symmetry. The steady state which is reached at large time is found to depend on some features of the initial conditions. By contrast, both of these features are absent in the case of Brownian particles, for which the Gibbs equilibrium is unique. We note that, although this work focuses on the large $N$ limit of the model, the existence of multiple steady states and the breaking of parity symmetry are already visible for small values of $N$, as we have observed from numerical simulations, which we report in Appendix B.
Although the above results are obtained for an analytically tractable “toy model”, i.e. a double well interaction potential which grows at infinity, one can make an argument that a similar transition occurs for more realistic interaction forces which vanish at infinity, but for which the system still has a bound state. A more precise criterion for the existence of a bound state and for the transition from single to disjoint support is discussed in Appendix C, and is confirmed by a numerical simulation. In that case we expect all the results obtained for the double well “toy model” to be qualitatively relevant.
@@ I.3 Main results
In order to test these analytical predictions, obtained in the limit $N\to+\infty$, we have performed direct numerical simulations of the equation of motion (1). The stationary density is determined by averaging over a large time window. We find a good agreement between theory and numerics already for $N=100$ (see Fig. 3). In particular, both the bistability (see Fig. 5) and the existence of asymmetric steady-states (see Fig. 8) are confirmed by these numerical results. The case of small values of $N$ is discussed in Appendix B.
@@ II.2 Calculation of the stationary density in the purely active case
Finally, we have tested these predictions using numerical simulations. We solve numerically the equation of motion (1) at $T=0$ for $N=100$ particles, starting from a uniform initial condition on some interval $x_{init},x_{init}]$ and we determine the stationary total density by time averaging the density in the center of mass frame at large time. For $k>k_{c}$, we instead start from an initial condition where the particles are separated into two groups of equal size, which allows to obtain a symmetric stationary state. Different choices of initial condition may lead to an asymmetric stationary state, see Sec. III. The results are displayed in Fig. 3 and show excellent agreement.
Figure 3: Plots of the density $\rho_{s}(x)$ for different values of $k$ (corresponding to different regimes in the diagram of Fig. 1 and for $\gamma=2$ (left) and $\gamma=6$ (right) (with $g=1$, $v_{0}=1$ and $x_{init}=1$). The plots in light colors show the theoretical prediction (given by (29) or (43)) while the darker lines were obtained by simulating the stochastic dynamics for $N=100$ particles and averaging the histogram of positions in the steady state. For each value of $k$, the corresponding value of $k_{0}$ to be used in the simulations was computed numerically using the relation (11).
@@ II.3.1 Relation between $k$ and $k_{0}$
Figure 4: Top left: Plot of $m_{2}$ versus $k$ for different values of $\gamma$. In the limit $\gamma\to 0^{+}$ (black curve), $m_{2}(k)$ has a simple expression given in (47). In this case, the function $m_{2}(k)$ is discontinuous at $k_{c}$, but it is smooth for any $\gamma>0$. Top right: Same plot zoomed around $k_{c}$, and for values of $\gamma$ close to $\gamma_{c}$. While there seems to be a cusp at $k_{c}$ for small $\gamma>0$ when looking at the large scales, the function $m_{2}(k)$ appears to be smooth when zooming sufficiently. Bottom left: Plot of $k$ versus $k_{0}$. When $\gamma$ is smaller than some critical value $\gamma_{c}=0.1787369$..., the function $k_{0}(k)$ becomes non-monotonous close to $k_{c}$. The inverse function $k(k_{0})$ thus becomes multi-valued, leading to the coexistence of 2 stable and one unstable steady states for the same value of $k_{0}$, corresponding to different values of $k$. The value $\gamma_{c}=0.1787369$... was computed numerically, using as a criterion that the minimum of $\frac{dk_{0}}{dk}$ vs $k$ vanishes at $\gamma_{c}$ (it is negative for $\gamma<\gamma_{c}$ and positive for $\gamma>\gamma_{c}$). For $\gamma=0^{+}$, the bistability occurs in the interval $k_{0}\in[6.614585...,9.449407$...]. Bottom right: Same plot zoomed around $k_{c}$, and for values of $\gamma$ close to $\gamma_{c}$. The region where the function $k_{0}(k)$ is non-monotonous seems to be entirely located at $k<k_{c}$.
@@ II.3.3 Bistability
Figure 5: Example where two steady states are observed in simulations for the same values of the parameters ($k_{0}=8$, $\gamma=0.01$, $g=1$, $v_{0}=1$ and $N=100$). The two densities were obtained using the same uniform initial condition with support [1,1][-1,1]. One of steady states (in blue) has a single support, while the other one (in orange) has a disjoint support. For each of the two steady states, the measured values of $m_{2}$ and $k$ ($k_{em}$ in the legend of the figure) coincide quite well with the two possible values predicted theoretically ($k_{\rm th}$ in the legend). The lighter lines show the two theoretical predictions for $N\to+\infty$ for the two possible values of $k$, which for such a small value of $\gamma$ are very to delta peaks at $x=\pm y_{1}$, and at $x=\pm y_{3}$ in the case of the disjoint support.
For $\gamma<\gamma_{c}$, there exists an interval of values of $k_{0}$ around $k_{0,c}$ where there are three possible values of kk [^1] Note that for a given value of $k$ there is however a unique corresponding value of $k_{0}$, hence in particular the critical value $k_{0,c}=k_{0}(k_{c})$ is uniquely defined., see Fig. 4. The intermediate value, which is always very close to $k_{c}$, seems to be unstable and was never observed in the simulations. The other two values however can both be realized, depending on the initial condition and on the realization of the noise, leading to two possible stationary states. In most cases one of these stationary state has a connected support ($k<k_{c}$) and the other one has a disconnected support ($k>k_{c}$). An example is shown in Fig. 5 (for $N=100$ particles), where either one or the other stationary state is reached depending on the realization, starting from the same initial condition. In a small region of parameters $k_{0}$ and $\gamma$ however, we have noticed that both stationary states correspond to $k<k_{c}$, i.e. to a connected support, see Fig. 4. An important consequence of this bistability is that in some range of parameters the transition between a connected and disconnected support is now discontinuous as a function of $k_{0}$.
Finally, there are two interesting questions for future investigations. It would be nice to be able to prove the stability properties of the different stationary solutions at infinite $N$. It is an open question whether at finite $N$ the system can transition from one stationary state to the other over an extremely long period of time.
@@ III Non symmetric steady state
Until now we have assumed that the stationary density $\rho_{s}(x)$ is symmetric, $\rho_{s}(-x)=\rho_{s}(x)$. This should be true in the large $N$ limit if we start from a symmetric initial condition. However, as we now show, if the initial condition is not symmetric, and if the support is disjoint, a non-symmetric stationary distribution is possible. Such a distribution is characterized by a non-zero value of the third moment $m_{3}$.
@@ III.1 General properties
with
$\tilde{F}(x)=-\int dy(-k_{0}(x-y)+(x-y)^{3})\rho_{s}(y)=m_{3}+(k_{0}-3m_{2})x-x^{3}\;.$ (61)
The idea is to solve the first equation (59) for $\rho_{s}(x)$ with some arbitrary given $F(x)$, and then insert the solution in the r.h.s. of the second equation (60), which leads to a self consistent condition for $F(x)$, hence for $\rho_{s}(x)$. In practice we see that we can choose $F(x)=\mu_{3}+kx-x^{3}$, where $k$ and $\mu_{3}$ must be determined by (60). Equation (61) then gives the condition $k=k_{0}-3m_{2}$, as in the symmetric case, as well as $\mu_{3}=m_{3}$, where $m_{2}$ and $m_{3}$ should be interpreted as functions of $k$ and $\mu_{3}$.
@@ IV Conclusion
It is quite interesting that this simple many body system exhibits a non-uniqueness of the stationary state in one space dimension. For the case of thermal Brownian particles there is a unique Gibbs state at any finite temperature $T>0$, while symmetry breaking and non-uniqueness only occur at $T=0$. The much richer phenomena observed here seem to be a consequence of the boundedness of the telegraphic RTP noise. It would be very interesting to investigate whether such phenomena occur for similar interactions with other kind of active particles and in higher space dimensions.

### v3

@@ I.1 Overview
In this paper we consider a toy model for such interactions, which consists in $N$ one-dimensional RTPs interacting via a pairwise polynomial potential of the form $W(r)=-k_{0}\,r^{2}/2+g\,r^{4}/4$, which for $k_{0},g>0$ is repulsive at short distance and attractive at large distance. We find that in the steady state the system forms a bound state with a bounded support, which exhibits a transition between a joint support and a disconnected support where particles spontaneously split into two groups. Furthermore, remarkably, we find that the steady state is not unique, and that in some range of parameters it exhibits bistability, as well as a breaking of the parity symmetry. The steady state which is reached at large time is found to depend on some features of the initial conditions. By contrast, both of these features are absent in the case of Brownian particles, for which the Gibbs equilibrium is unique. We note that, although this work focuses on the large $N$ limit of the model, the existence of multiple steady states and the breaking of parity symmetry are already visible for small values of $N$, as we have observed from numerical simulations, which we report in Appendix B.
Our model has two advantages from a computational perspective. First, $W^{\prime}(0)=0$, hence it allows for particle crossings and therefore allows to use the extension of the Dean-Kawasaki method Dean ; Kawa to the RTP (while it was shown that this method fails to describe single-file active particles TouzoDBM2023 ). Second, it is simple enough so that the self-consistent equation obeyed by the stationary density in the large $N$ limit, derived in activeRDshort , can be solved analytically. Furthermore, although our model has an interaction force $W^{\prime}(r)$ which diverges at large distance, which is a priori an unrealistic feature, one can make an argument that a similar behavior can be observed for more realistic interaction forces which vanish at infinity, but for which the system still has a bound state. A more precise criterion for the existence of a bound state and for the transition from single to disjoint support is discussed in Appendix C, and is confirmed by a numerical simulation. In that case we expect all the results obtained for the double well toy model to be qualitatively relevant.
@@ I.3 Main results
Finally, a limit of interest for RTPs is the diffusive limit obtained by taking $\gamma,v_{0}\to+\infty$ with $T_{\rm eff}=\frac{v_{0}^{2}}{2\gamma}$ fixed (at fixed $k_{0}$). In that limit one expects to recover the results for Brownian particles. One can see from (8) that it corresponds to $\tilde{k}_{0}\to 0$, hence to a support of the density which is always joint, as expected from Gibbs equilibrium at temperature $T_{\rm eff}$.
In order to test these analytical predictions, obtained in the limit $N\to+\infty$, we have performed direct numerical simulations of the equation of motion (1). The stationary density is determined by averaging over a large time window. We find a good agreement between theory and numerics already for $N=100$ (see Fig. 3). In particular, both the bistability (see Fig. 5) and the existence of asymmetric steady states (see Fig. 8) are confirmed by these numerical results. The case of small values of $N$ is discussed in Appendix B.
@@ II.2 Calculation of the stationary density in the purely active case
Finally, we have tested these predictions using numerical simulations. We solve numerically the equation of motion (1) at $T=0$ for $N=100$ particles, starting from a uniform initial condition on some interval $x_{init},x_{init}]$ and we determine the stationary total density by time averaging the density in the center of mass frame at large time. For $k>k_{c}$, we instead start from an initial condition where the particles are separated into two groups of equal size, which allows to obtain a symmetric stationary state. Different choices of initial condition may lead to an asymmetric stationary state, see Sec. III. The results are displayed in Fig. 3 and show excellent agreement (although finite $N$ effects are visible near the edges of the support in some cases).
Figure 3: Plots of the density $\rho_{s}(x)$ for different values of $k$ (corresponding to different regimes in the diagram of Fig. 1 and for $\gamma=2$ (left) and $\gamma=6$ (right) (with $g=1$, $v_{0}=1$ and $x_{init}=1$). The plots in light colors show the theoretical prediction (given by (29) or (43)) while the darker lines were obtained by simulating the stochastic dynamics for $N=100$ particles and averaging the histogram of positions in the steady state. For each value of $k$, the corresponding value of $k_{0}$ to be used in the simulations was computed numerically using the relation (11), and is given in Appendix D.
@@ II.3.1 Relation between $k$ and $k_{0}$
Figure 4: Top left: Plot of $m_{2}$ versus $k$ for different values of $\gamma$, obtained by evaluating numerically the integral in (31) for $k>k_{c}$, and from an equivalent expression deduced from (43) for $k<k_{c}$. In the limit $\gamma\to 0^{+}$ (black curve), $m_{2}(k)$ has a simple expression given in (47). In this case, the function $m_{2}(k)$ is discontinuous at $k_{c}$, but it is smooth for any $\gamma>0$. Top right: Same plot zoomed around $k_{c}$, and for values of $\gamma$ close to $\gamma_{c}$. While there seems to be a cusp at $k_{c}$ for small $\gamma>0$ when looking at the large scales, the function $m_{2}(k)$ appears to be smooth when zooming sufficiently. Bottom left: Plot of $k$ versus $k_{0}$ (obtained by the same method). When $\gamma$ is smaller than some critical value $\gamma_{c}=0.1787369$..., the function $k_{0}(k)$ becomes non-monotonous close to $k_{c}$. The inverse function $k(k_{0})$ thus becomes multi-valued, leading to the coexistence of 2 stable and one unstable steady states for the same value of $k_{0}$, corresponding to different values of $k$. The value $\gamma_{c}=0.1787369$... was computed numerically, using as a criterion that the minimum of $\frac{dk_{0}}{dk}$ vs $k$ vanishes at $\gamma_{c}$ (it is negative for $\gamma<\gamma_{c}$ and positive for $\gamma>\gamma_{c}$). For $\gamma=0^{+}$, the bistability occurs in the interval $k_{0}\in[6.614585...,9.449407$...]. Bottom right: Same plot zoomed around $k_{c}$, and for values of $\gamma$ close to $\gamma_{c}$. The region where the function $k_{0}(k)$ is non-monotonous seems to be entirely located at $k<k_{c}$.
@@ II.3.3 Bistability
Figure 5: Example where two steady states are observed in simulations for the same values of the parameters ($k_{0}=8$, $\gamma=0.01$, $g=1$, $v_{0}=1$ and $N=100$). The two densities were obtained using the same uniform initial condition with support [1,1][-1,1], but for different realizations of the telegraphic noise. One of steady states (in blue) has a single support, while the other one (in orange) has a disjoint support. For each of the two steady states, the measured values of $m_{2}$ and $k$ ($k_{em}$ in the legend of the figure) coincide quite well with the two possible values predicted theoretically ($k_{\rm th}$ in the legend). The lighter lines show the two theoretical predictions for $N\to+\infty$ for the two possible values of $k$, which for such a small value of $\gamma$ are very close to delta peaks at $x=\pm y_{1}$, and at $x=\pm y_{3}$ in the case of the disjoint support.
For $N\to+\infty$ and $\gamma<\gamma_{c}$, there exists an interval of values of $k_{0}$ around $k_{0,c}$ where there are three possible values of kk [^1] Note that for a given value of $k$ there is however a unique corresponding value of $k_{0}$, hence in particular the critical value $k_{0,c}=k_{0}(k_{c})$ is uniquely defined., see Fig. 4 (bottom). There are two possible situations. (i) For most values of $\gamma$ and $k_{0}$, two of these stationary state have a connected support ($k<k_{c}$) and the other one has a disconnected support ($k>k_{c}$). (ii) In a small region of parameters $k_{0}$ and $\gamma$ however, we have noticed that all three solutions correspond to $k<k_{c}$, i.e. to a connected support, see Fig. 4 (bottom right). All three values correspond a priori to possible stationary states of the system, which however can be stable or unstable. We have not studied analytically the stability of these solutions (which seems a priori technically challenging). However, we have compared these predictions with numerical simulations for finite $N$ as we now discuss.
The intermediate value, which is always very close to $k_{c}$, seems to be unstable and was never observed in the simulations. In case (i), the other two values however can both be realized, depending on the initial condition and on the realization of the noise, leading to two possible stationary states. An example from numerical simulations is shown in Fig. 5 (for $N=100$ particles), where either one or the other stationary state is reached depending on the realization, starting from the same initial condition. The case (ii) is more difficult to observe in numerical simulations as it corresponds to a very limited range of parameters, and it is not clear whether both solutions are indeed stable in this case.
Note that an important consequence of this bistability is that in some range of parameters the transition between a connected and disconnected support is now discontinuous as a function of $k_{0}$. Finally, there are two interesting questions for future investigations. It would be nice to be able to prove the stability properties of the different stationary solutions at infinite $N$. It is an open question whether at finite $N$ the system can transition from one stationary state to the other over an extremely long period of time.
@@ III Non symmetric steady state
Until now we have assumed that the stationary density $\rho_{s}(x)$ is symmetric, $\rho_{s}(-x)=\rho_{s}(x)$. This should be true in the large $N$ limit if we start from a symmetric initial condition. It should also hold whenever the support has a single component, due to ergodicity. However, as we now show, if the initial condition is not symmetric, and if the support is disjoint, a non-symmetric stationary distribution is possible, where the particles are separated into two groups of different sizes. Such a distribution is characterized by a non-zero value of the third moment $m_{3}$.
@@ III.1 General properties
where
$\tilde{F}(x)=-\int dy(-k_{0}(x-y)+(x-y)^{3})\rho_{s}(y)=m_{3}+(k_{0}-3m_{2})x-x^{3}\;,$ (61)
is the force that is obtained from the steady state density. The idea is to solve the first equation (59) for $\rho_{s}(x)$ with some arbitrary given $F(x)$, and then insert the solution in the r.h.s. of the second equation (60), which leads to a self consistent condition for $F(x)$, hence for $\rho_{s}(x)$. In practice we see that we can choose $F(x)=\mu_{3}+kx-x^{3}$, where $k$ and $\mu_{3}$ must be determined by (60). Equation (61) then gives the condition $k=k_{0}-3m_{2}$, as in the symmetric case, as well as $\mu_{3}=m_{3}$, where $m_{2}$ and $m_{3}$ should be interpreted as functions of $k$ and $\mu_{3}$.
@@ IV Conclusion
It is quite interesting that this simple many body system exhibits a non-uniqueness of the stationary state in one space dimension. For the case of thermal Brownian particles there is a unique Gibbs state at any finite temperature $T>0$, while symmetry breaking and non-uniqueness only occur at $T=0$. The much richer phenomena observed here seem to be a consequence of the boundedness of the telegraphic RTP noise. In particular, we expect that adding a small thermal noise to the model would lead to an unbounded support, and transform the non-symmetric steady-states into long lived transients towards a symmetric steady state. It is however less clear what would happen to the bistability discussed in Sec. II.3.3. Understanding more generally the conditions on the noise and the interaction to observe such non-uniqueness of the stationary state is an open direction for future research. It would be particularly interesting to investigate whether such phenomena occur for similar interactions with other kinds of active particles and in higher space dimensions.
@@ C.2 Limit $\gamma\to 0$
Appendix D Values of $k_{0}$ for Fig. 3
$k$ -6.5 -3 -1 1 2 3 4
$\gamma=2$ -2.862008 -0.556904 3.561302 7.783972 11.895498
$\gamma=6$ -6.475127 -0.803101 3.703725 7.900638 11.943203 15.962077
Table 1: Correspondence between the values of $k$ used for the analytical predictions and the values of $k_{0}$ used for the numerical predictions on the two plots of Fig. 3, computed using the relation (11).